A tax system has the following income tax brackets: $0 to $20,000, 10%; $20,001 to $40,000, 15%; $40,001 to $60,000, 20%; and over $60,000, 25%. Calculate the effective tax rate and marginal tax rate for both of these individuals: Tom makes $90,000 and has $10,000 in itemized deductions and $3,500 in a personal exemption. John makes $35,000 and has $5,900 in the standard deduction and a $3,500 personal exemption. Based on your calculations for these individuals, is the tax system regressive, progressive or proportional? Why?
Answer #1

Tom’s taxable income

Adjusted gross income

90000

Less: itemized deductions

(10000)

Less: personal exemption

(3500)

Taxable income

$76500

Tax liability (20000*10%)+(15%*(40000-20000))+(20%*(60000-40000))+(25%*(76500-60000))

$13125

Effective tax rate = tax liability / total income = 13125/90000 = 14.58%

Marginal tax rate = 25%

John’s taxable income

Adjusted gross income

35000

Less: itemized deductions

(5900)

Less: personal exemption

(3500)

Taxable income

$25600

Tax liability (20000*10%)+(15%*(25600-20000))

$2840

Effective tax rate = tax liability / total income = 2840/35000 = 8.11%

Marginal tax rate = 15%

The tax system is progressive because the tax rate increase with the increase in income. However the increase in tax rate is more than the proportional increase in income.